OpenStudy (anonymous):

How do you find the ingral of= ʃ(Chx)^2 dx

zepdrix (zepdrix):

What are C and h? Constants?

OpenStudy (anonymous):

CHx it's a hyperbolic function.

OpenStudy (turingtest):

you mean\[\int \cosh^2 xdx\]?

OpenStudy (turingtest):

hyperbolic cosine function, yes?

OpenStudy (anonymous):

\[\int\limits_{}^{} (Chx )^{2} dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Chx = Coshx = hyperbolic cosine function

OpenStudy (turingtest):

use the definition\[\cosh x={e^x+e^{-x}\over2}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } \int\limits_{}^{} (e ^{x} + e ^{-x})^{2} dx\]

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

\[\frac{1}{4}\int\limits_{}^{} (e ^{2x} + e ^{-2x} + 2 ) dx\]

OpenStudy (anonymous):

Then how should I proceed?

OpenStudy (anonymous):

Integrate term-by-term.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @naylah \[\frac{1}{4}\int\limits_{}^{} (e ^{2x} + e ^{-2x} + 2 ) dx=\frac{1}{4}\int e ^{2x}~dx + \frac{1}{4}\int e ^{-2x}~dx + \frac{1}{2}\int dx\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } (\frac{ e ^{2x} }{ 2 }) + \frac{ 1 }{ 4 } (\frac{ e ^{-2x} }{ -2 }) + \frac{ x }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } (\frac{ e ^{2x} - e ^{-2x} }{ 4 }) + \frac{ x }{ 2 } + c\]

OpenStudy (anonymous):

\[\frac{ (Shx) ^{2} }{ 2 } + \frac{ x }{ 2 } + c\]

OpenStudy (anonymous):

I did as ypu told me but on the book the result is different. On the book it says that the result is: \[\frac{ (ShxChx + x) }{ 2 } + c\]

zepdrix (zepdrix):

\[\Large\bf\sf \left(\frac{e^{2x}-e^{-2x}}{4}\right)\quad\ne\quad \frac{(Shx)^2}{2}\] \[\Large\bf\sf (Shx)^2\quad=\quad \left(\frac{e^x-e^{-x}}{2}\right)^2\quad=\quad \frac{1}{4}\left(e^{2x}-2+e^{-2x}\right)\] We would need that extra middle term to simplify it down to Shx, yes?

zepdrix (zepdrix):

Oh I guess it might be better to do something like this.\[\Large\bf\sf \frac{1}{4}\left(\frac{e^{2x}-e^{-2x}}{2}\right)\quad=\quad \frac{1}{4}\sinh(2x)\quad=\quad \frac{1}{2}Shx \cdot Chx\]Applying Double Angle Formula.

OpenStudy (anonymous):

Sorry for my mistake and thank you :D